The perimeter of a rectangle is 90 m . Its length is 1 & a half times its width. Find the length & width of the rectangle.
Answer by richwmiller(17219)

let W=width
let L=length
L=1.5w
L+L+W+W=90 meters
1.5W +1.5W +2w=90
5W=90
width=W=18 meters
l=1.5 w=1.5(18)
length= L=18+9 =27 meters
check
27+27+18+18=90
54+36=90
90 =90